(y-7)^2=16(x+5)
how do I make this a y= equation for a parabola?

Answer by stanbon(75887)

(y-7)^2=16(x+5)
----
You don't.
It's an "x=" parabola.
---
x = (1/16)(y-7)^2-5
---
It opens to the right from vertex (7,-5)
=======================
Cheers,
Stan H.